OpenStudy (anonymous):

I need someone who can help me understand how to solve Series.I have an exam tommorow,so any help would be much appreciated :)

OpenStudy (anonymous):

\[\sum_{\infty}^{1}\frac{ n }{ n ^{2} +1}\]

OpenStudy (anonymous):

Thank you for calling backup :D

OpenStudy (anonymous):

@ayeshaafzal221

OpenStudy (anonymous):

do u want to learn basic of series ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok :) give me one minute

OpenStudy (anonymous):

Sure ^^

OpenStudy (anonymous):

ok so there is two types of series one is infinite which is set of positive numbers and keeps going on theres no stop to it a1, a2 , a3 ,...... then second type is finite which means limited amount of numbers are in that sequence.

OpenStudy (anonymous):

I have to learn the infinite ones

OpenStudy (anonymous):

Like the one I posted above

OpenStudy (anonymous):

ok can u show me some example of what kind of question you will be getting so i want to know which level u are at , if you dont mind.

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

I choose the ones that looked easier for me to understand and put them in that .jpg

OpenStudy (anonymous):

I would like to solve them one by one so I can understand them

OpenStudy (anonymous):

u know how to take limits right?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

i just learned the basic with another nice guy from this site :D

OpenStudy (anonymous):

ok its quite easy basically you take the limit and divide each variable of question with biggest value it ll be mor clear once i solve the question

OpenStudy (anonymous):

\[\sum_{\infty}^{1}\frac{ n }{ n ^{2}+1 }\]

OpenStudy (anonymous):

i divide with n^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but take the limit first like \[\lim_{n \rightarrow \infty} \frac{ n }{ n ^{2} +1}\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty }\frac{ \frac{ n }{n ^{2}} }{ \frac{ n ^{2} }{ n ^{2} } +1}\]

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

And now?

OpenStudy (anonymous):

now you cross out the like terms

OpenStudy (irishboy123):

this might help out http://www.math.hawaii.edu/~ralph/Classes/242/SeriesConvTests.pdf

OpenStudy (anonymous):

If I cross out the like terms,wouldn't it lead me to the exact ecuation with which we started?

OpenStudy (anonymous):

no it will be \[\frac{ 0 }{ 1+1 } =0\]

OpenStudy (anonymous):

Ohh,I see

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

\[\frac{ 0 }{ 1+0 } =0\]

OpenStudy (anonymous):

it is infinite/infinite ^ 2 = 0 because infinite ^ 2 is bigger

OpenStudy (anonymous):

and infinite ^ 2 / infinite ^ 2 = 1 because it is the same?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do u know what convergent and divergent means?

OpenStudy (anonymous):

something that has to do with the result being higher or lower than 1

OpenStudy (anonymous):

yes now can u go to this website it has everything you need to know , go through and message me if you dont understand a thing its with examples and solution

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

dont get scared of first thing scroll down

OpenStudy (anonymous):

your welcome :0

OpenStudy (anonymous):

:)*

OpenStudy (anonymous):

google khan academy geometric series , its a life saver

OpenStudy (anonymous):

make a free login

OpenStudy (anonymous):

I will try to start solve the exercises to see if I can,when I''m stuck I will write the exercise here

OpenStudy (anonymous):

Jhanny

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

\[\sum_{\infty}^{1}\frac{ n^{2} }{2^{n} }\]

OpenStudy (anonymous):

anyone can help me solve this?

OpenStudy (irishboy123):

put it in a new thread

OpenStudy (anonymous):

ok